# The Design Space of Lockfiles Across Package Managers

### v2

@@ Abstract
We perform the first comprehensive study of lockfiles across 7 popular package managers, npm, pnpm, Cargo, Poetry, Pipenv, Gradle, and Go. First, we highlight the wide variety of design decisions that package managers make, regarding the generation process as well as the content of lockfiles. Next, we conduct a qualitative analysis based on semi-structured interviews with 15 developers. We capture first-hand insights about the benefits that developers perceive in lockfiles, as well as the challenges they face to manage these files. Following these observations, we make 5 recommendations to further improve lockfiles, for a better developer experience.
@@ 1 Introduction
Software reuse is an undeniable best practice in software engineering [1]. Software developers reuse third-party libraries for commodity features such as logging, cryptography, arithmetic, or I/O. Reuse helps developers save time, pool quality assurance efforts, and focus on the core, application specific features of their project [2]. Yet, software reuse at scale introduces major engineering challenges [3]. The use of third-party libraries, including the transitive ones, often results in large dependency trees. Managing these increasingly complex dependency trees is hard because developers have to resolve compatible versions across the whole tree [4], stay on top of updates to patch bugs and vulnerabilities [5], and ensure that new changes do not break the build [6].
@@
Package managers are tools that assist developers in the management of reused libraries [7]. They automate and simplify the configuration, resolution, installation, and update of dependencies. When using a package manager, developers define the dependencies of a project in a dependency specification file. Then, the package manager determines compatible versions of these dependencies and all indirect dependencies, fetches them from a package registry, and installs them to further build the project. During this process, some package managers save the exact list of resolved dependencies and their versions into a special file: the lockfile [8].
@@
Lockfiles are widely used throughout the software development lifecycle, supporting reproducible builds [9], defending the software supply chain [10], and contributing to improved build performance [11]. However, lockfiles also introduce additional complexity, and their maintenance requirements are not fully understood, a subject hardly touched by software engineering research. Lockfiles have been mentioned as security best practices [12], important for software supply chain security [13], related to Software Bill of Materials (SBOM) generation [14, 8], and notable for new dependency management tools [15]. However, none of these works provide an in-depth analysis of lockfiles, why they exist and what they should contain. To our knowledge, there is no systematic analysis and comparison of the different strategies in the field to generate and evolve lockfiles. There is no grounded evidence on how developers perceive the benefits and challenges of maintaining lockfiles.
@@
In this study, we address this critical gap in the software engineering literature by conducting a comprehensive analysis of lockfiles across 7 package managers. These package managers serve the top programming languages of 2024 [16]: npm and pnpm for Javascript, Cargo for Rust, Poetry and Pipenv for Python, Gradle for Java, and Go for Golang. We analyze the implementation of these package managers, to unveil the large design space for the generation, maintenance and content of lockfiles. We then quantitatively evaluate lockfile usage in the wild, by collecting GitHub projects and their lockfile commit rates. To complement this analysis, we conduct interviews with 15 developers with diverse backgrounds and experiences. They contribute to active open source projects spanning various domains such as networking, gaming, CLI, and APIs, and use the package managers under study. Through these interviews, we gather valuable and grounded first-hand insights into the developers’ motivations for using or avoiding lockfiles, as well as the challenges they face when working with them.
@@
Our paper is the first study that highlights how the functionalities and usage of lockfiles differ across package managers and ecosystems, resulting in varying levels of developer satisfaction. While some approaches, such as the Go strategy for lockfiles, encourage strong developer engagement, others, such as Gradle’s approach, fail to address key software engineering challenges. Our findings clearly show that a simple, principled approach to the lockfile lifecycle leads to a better developer experience. We also gather evidence of the challenges that developers face with lockfiles, such as limited human readability, difficulty locking indirect dependencies, and delays in receiving dependency updates. Overall, developers generally trust lockfiles and view them as a valuable tool for improving the efficiency and security of dependency management. We conclude our study with 5 recommendations to encourage the adoption of lockfiles in all projects. Our findings are directly actionable by package manager users and by package manager maintainers. Our results assist package manager users in understanding lockfiles and choosing a package manager that better suits their project requirements. Our findings help package manager maintainers improve their lockfile feature, which in turn can positively impact a large community of developers. All data and interview protocols used in this study are publicly available: https://github.com/chains-project/lockfiles-comprehensive-study. In summary, our key contributions are,
@@
  • the first comprehensive comparison of lockfiles across the 7 top package managers, with a deep analysis of the key dimensions of the lockfile problem space: the lockfile lifecycle, and content of the generated lockfile
@@
  • a regimen of 2 recommendations for package manager users and 3 recommendations for package manager developers to further improve lockfile usage for dependency management, to inform future research in the field as well as engineering of the future generation of package managers.
@@ 2.1 Definitions
Figure 1: A canonical workflow of interactions between a developer, the dependency specification file, the package manager, the lockfile, and the package registry.
@@
A package manager facilitates code reuse by ensuring that all the third-party packages required by the project, directly or indirectly, are downloaded with the correct versions. Most popular programming languages have one or more package managers (e.g. Maven for Java or Cargo for Rust). Package managers typically provide developers with a command line interface and sometimes a graphical user interface to interact with the package registry.
@@
A package registry is as a central platform for sharing code across multiple projects. Depending on the registry, developers can upload packaged binaries (e.g., JAR files of Java bytecode) or packaged source code, allowing others to download the code to be reused within their own projects. Popular programming languages have official package registries: npm.com is widely used for JavaScript, Crates.io for Rust, PyPI.org for Python, etc.
@@
A dependency specification is a file that specifies the packages that a software project directly depends on, i.e., it specifies the direct dependencies of the project. For each dependency, the developers can either declare a specific version, a practice commonly known as pinning, or specify a range of versions, or specify no version which usually means allow resolving the ‘latest compatible‘ version. The package manager takes the dependency specification file as input to build the project. The dependency specification file also includes additional information related to dependency management, such as metadata about the project or exclusion of certain dependencies. The dependency specification file is primarily maintained by developers, and also by automated dependency bots such as Dependabot.
@@
A lockfile is a file that is automatically generated by the package manager and that is not meant to be edited by developers. The lockfile records the resolved versions of all direct and indirect dependencies that need to be included in a project. Some package managers may include extra details, such as the hash of the dependency specification file content or the package manager version. Package managers use lockfiles when (re)building projects, such as for fetching the exact versions of dependencies required, and verifying their integrity. Lockfiles may also be read by developers during code reviews to audit resolved dependencies, and updated by automated dependency bots during dependency updates.
Table 1: The 7 package managers analyzed in this study, including their associated programming languages, registries, dependency specifications, and lockfiles.
@@ 2.2 Considered Package Managers
We select 7 popular package managers for our study based on the popularity of programming languages in 2024 [16]. From the ten most popular languages, we identify those with widely used package managers that support lockfiles: JavaScript, TypeScript, Python, Java, and Go. We add Rust due to its significant upward trend. For each language, we select the most commonly used package manager with lockfile support: npm for JavaScript/TypeScript, Poetry for Python, Cargo for Rust, Gradle for Java, and Go’s built-in toolchain. When a language has multiple package managers, we add a second one: pnpm for JavaScript and Pipenv for Python. Table 1 outlines the 7 package managers analyzed in this study, along with their associated languages, registries, dependency specification files, and lockfiles.
@@
For instance, npm CLI is the default package manager for JavaScript, and the npm registry is the default package registry. Library developers can upload their JavaScript code to the npmjs.com registry. The dependency specification file and lockfile for npm CLI are in JSON format and are named package.json and package-lock.json, respectively. Excerpts from a package.json and its corresponding package-lock.json are given in 1 and 2, respectively.
@@
In the case of Go, there is no centralized package registry. Instead, Go modules are downloaded from version control registries such as GitHub. Module versions are tracked using an index that feeds available new versions. The integrity verification of Go packages is done based on a checksum database. Moreover, in Go, the go.mod file works both as the dependency specification file and the lockfile. An additional file, named go.sum, is generated by the package manager, to record the checksums of the downloaded modules. The go.mod file is not meant to be edited manually by developers but should be modified using specific commands. 3 and 4 show an example go.mod file and a go.sum file.
@@
Table 2 provides the list of selected package manager versions, including the number of lines of code in the specified version, and the number of available releases up to that version. The release counts are obtained from the official releases page of each package manager, except for pnpm, where the release count is based on the information from the npm registry.
@@
Table 2: Details about the exact versions of the 7 package managers under study.
@@
Package manager Analyzed Version Lines of code Versions
npm CLI npm/cli (11.1.0) 459k (JavaScript) 557
pnpm pnpm/pnpm (10.1.0) 112k (TypeScript) 1103
Cargo rust-lang/cargo (0.85.0 ) 210k (Rust) 119
Poetry python-poetry/poetry (2.0.1) 46k (Python) 88
Pipenv pypa/pipenv (2024.4.1) 140k (Python) 396
Gradle gradle/gradle (8.12.1) 784k (Groovy) 531k (Java) 367k (Kotlin) 148
Go golang/go (1.24.0) 2052k (Go) 243
@@ 2. RQ2:
    To what extent do open-source projects include lockfiles in version control?
    In this RQ, we quantitatively evaluate how developers of different ecosystems include lockfiles in version control. Including lockfiles in version control systems is important to achieve all benefits lockfiles have to offer, including deterministic builds [17]. We base our evaluation on lockfile commit data collected from open source GitHub projects created within the past 5 years (from 2019 to 2024). This quantitative analysis helps consolidate our observations in RQ1 about the usability and usefulness of lockfiles.
@@ 3.2 Methodology for RQ1
To answer the RQ1, we analyze the implementations of the lockfile feature in the source code of package managers, we review their documentation, and we experiment with sample projects that use each package manager.
Source code analysis: To understand the workflows of lockfile-related tasks for each package manager, we analyze their implementation within their source code repository. We begin by locating the main function that is invoked after executing an install or build command. We then analyze the functions called by this main function, manually tracing their execution flow. We further search the project using regular expressions for the terms: lockfile, locked, frozen, and dependency resolution. The sections of source code, analyzed for each package manager are summarized in Table 3. For instance, in the case of npm, we study the repository npm/cli, focusing on the source code for install and ci actions, dependency tree resolution, and lockfile validation.
Documentation: In addition to analyzing the source code of the 7 package managers, we review their official documentation. We focus on sections related to dependency resolution actions such as install or build, and study the available configurations or flags that can be used with these actions. We search for sections covering lockfiles, including lockfile-related recommendations and comparisons between lockfiles and dependency specification files. For example, as given in Table 3, for Poetry, we go through the section installing dependencies that explain lockfile usage, including installing without poetry.lock, installing with poetry.lock, and committing poetry.lock to version control. Similarly, for Pipenv, we analyze the commands and Pipfile & Pipfile.lock sections that cover general lockfile recommendations, examples, and security features.
Table 3: Lockfile features for the 7 package managers under study. We study their source code and documentations.
@@
3.3 Methodology for RQ2
Next we aim to analyze lockfile usage across package managers, by selecting a sample of open source projects and collecting the lockfiles within them, if any.
Project selection: As the study subjects for RQ2, we collect popular open-source GitHub projects that use one of the 7 package managers we consider in the study. When collecting GitHub projects, we aim to exclude toy projects and only target mature, widely-used projects. To achieve this, we only consider repositories created between September 30, 2019, and September 30, 2024, that have at least 42 stars, 10 contributors, and 300 commits. To ensure that we only select actively maintained repositories, we further exclude the projects that have not been updated within the last 3 months. For npm CLI and pnpm, we select repositories where the primary language is JavaScript or TypeScript. For Pipenv and Poetry, we include Python projects. For Cargo, Gradle, and Go, we target Rust, Java, and Go projects, respectively.
Lockfile identification: Similar to identifying the dependency specification file, we determine the presence of a lockfile by searching for the relevant lockfile files in the project directory. If a project has multiple lockfiles, we discard the project to prevent inconsistencies in the results and their interpretation. For example, in the JavaScript ecosystem, if a project contains both package-lock.json and pnpm-lock.json, we exclude it from the evaluation. We further disregard projects with lockfiles from package managers outside the selected 7 package managers, such as yarn.lock from Yarn. For Go, we check for the existence of the go.sum file. We also collect the project creation date and the earliest lockfile commit date. Using this information, we derive the number of projects that began including lockfiles in version control within the first six months of project creation, as well as the total count of projects that currently have lockfiles.
@@ 3.4 Results for RQ1: How do package managers generate, update, and structure lockfiles?
In this RQ, we first look at the fundamental differences in the lockfile content among the 7 selected package managers introduced in Table 1. Next we discuss the key differences in lockfile lifecycles including lockfile generation, updates and enforcement.

### v3

@@ Abstract
We perform the first comprehensive study of lockfiles across seven popular package managers, npm, pnpm, Cargo, Poetry, Pipenv, Gradle, and Go. First, we highlight the wide variety of design decisions that package managers make, regarding the generation process as well as the content of lockfiles. Next, we conduct a qualitative analysis based on semi-structured interviews with 15 developers. We capture first-hand insights about the benefits that developers perceive in lockfiles, as well as the challenges they face to manage these files. Following these observations, we make five recommendations to further improve lockfiles, for a better developer experience.
@@ 1 Introduction
Software reuse is an established best practice in software engineering [16]. Software developers reuse third-party libraries for commodity features such as logging, cryptography, arithmetic, or I/O. Reuse helps developers save time, pool quality assurance efforts, and focus on the core, application specific features of their project [25]. Yet, software reuse at scale introduces major engineering challenges [10]. The use of third-party libraries, including the indirect ones, often results in large dependency trees. Managing these increasingly complex dependency trees is hard because developers have to resolve compatible versions across the whole tree [33], stay on top of updates to patch bugs and vulnerabilities [21], and ensure that new changes do not break the build [30].
@@
Package managers are tools that assist developers in the management of reused libraries [38]. They automate and simplify the configuration, resolution, installation, and update of dependencies. When using a package manager, developers specify the dependencies of a project in a dependency specification file. The package manager then resolves compatible versions based on a combination of developer-specified constraints. Dependency resolution typically defaults to the latest available version if no version is specified. Then, the package manager fetches resolved dependencies from a package registry, and installs them to further build the project. During this process, some package managers save the exact list of resolved dependencies and their versions into a special file: the lockfile [48].
@@
Lockfiles are widely used throughout the software development lifecycle, supporting reproducible builds [22], defending the software supply chain [20, 24], and contributing to improved build performance [3]. However, lockfiles also introduce additional complexity, and their maintenance requirements are not fully understood, a subject hardly touched by software engineering research. Lockfiles have been mentioned as security best practices [18], important for software supply chain security [42], related to Software Bill of Materials (SBOM) generation [7, 48], and notable for new dependency management tools [9]. However, none of these works provide an in-depth analysis of lockfiles, why they exist and what they should contain. To our knowledge, there is no systematic analysis and comparison of the different strategies in the field to generate and evolve lockfiles. There is no grounded evidence on how developers perceive the benefits and challenges of maintaining lockfiles.
@@
In this study, we address this critical gap in the software engineering literature by conducting a comprehensive analysis of lockfiles across seven package managers. These package managers serve the top programming languages of 2024 [8]: Javascript, Python, Java and Golang. The considered package managers are: npm and pnpm for Javascript, Cargo for Rust, Poetry and Pipenv for Python, Gradle for Java, and Go for Golang. We deeply analyze the implementation of these package managers to uncover how they generate, update, and structure lockfiles. We then quantitatively evaluate the extent to which developers include lockfiles in version control by collecting GitHub projects and analyzing their lockfile commit rates. To complement this analysis, we conduct interviews with 15 developers from diverse backgrounds and experiences, who actively contribute to open-source. Through these interviews, we gather valuable and grounded first-hand insights into the benefits developers perceive in using lockfiles, as well as the challenges they encounter when working with them.
@@
Our paper is the first study that highlights how the functionalities and usage of lockfiles differ across package managers and ecosystems, resulting in varying levels of developer satisfaction. While some approaches, such as Go’s strategy to always generate a lockfile by default, facilitates their adoption by developers, others, such as Gradle that requires specializing the build in order to have a lockfile, fails to make the lockfile usable. Our findings clearly show that a simple, principled approach to the lockfile lifecycle leads to a better developer experience. We also gather evidence of the challenges that developers face with lockfiles, such as limited human readability, difficulty locking indirect dependencies, and delays in receiving dependency updates. Overall, developers generally trust lockfiles and view them as a valuable tool for improving the efficiency and security of dependency management. We conclude our study with five recommendations to encourage the adoption of lockfiles in all projects. Our findings are directly actionable by package manager users and by package manager maintainers. Our results assist package manager users in understanding lockfiles and choosing a package manager that better suits their project requirements. Our findings help package manager maintainers improve their lockfile feature, which in turn can positively impact a large community of developers. In summary, our key contributions are,
@@
  • the first comprehensive comparison of lockfiles across the seven top package managers, with a deep analysis of the key dimensions of the lockfile problem space: the lockfile lifecycle, and content of the generated lockfile.
@@
  • a regimen of two recommendations for package manager users and three recommendations for package manager developers to further improve lockfile usage for dependency management, to inform future research in the field as well as engineering of the future generation of package managers.
@@ 2.1 Definitions
Figure 1: A canonical workflow of interactions between a developer, the dependency specification file, the package manager, the lockfile, and the package registry. The order of events may slightly vary depending on the exact command executed by the developer in step 2.
@@
A package manager is a tool that automates code reuse by ensuring that all third-party packages required by the project, directly or indirectly, are downloaded with the correct versions. Most popular programming languages have one or more package managers. For example, Cargo is the package manager for Rust. Package managers typically provide developers with a command line interface and sometimes a graphical user interface to interact with the package registry.
@@
A package registry is as a central platform for sharing code across multiple projects. Depending on the registry, developers can upload packaged binaries or packaged source code, allowing others to download the code to be reused within their own projects. Popular programming languages have official package registries. Crates.io is the package registry for Rust and developers can distribute their packaged Rust source code through a central registry at Crates.io.
@@
A dependency specification is a file that specifies the packages that a software project directly depends on, i.e., it specifies the direct dependencies of the project. For example, the dependency specification file for Cargo is a file called Cargo.toml. For each dependency, developers can either declare a specific version, a practice commonly known as pinning, specify a range of versions, which usually resolves to the latest version within that range, or specify no version, which means resolving to the available latest version. The package manager takes the dependency specification file as input to build the project. The dependency specification file also includes additional information related to dependency management, such as metadata about the project or exclusion of certain dependencies. The dependency specification file is primarily maintained by developers, and also by automated dependency bots such as Dependabot.
@@
A lockfile is a file that is automatically generated by the package manager and that is not meant to be edited by developers. Each package manager uses its own file naming convention for the lockfile. For example, in Cargo, the file is named Cargo.lock. The lockfile records the resolved versions of all direct and indirect dependencies that need to be included in a project. Some package managers may include extra details, such as the hash of the dependency specification file content or the package manager version. Package managers use lockfiles when (re)building projects, such as for fetching the exact versions of dependencies required, and verifying their integrity. Lockfiles may also be read by developers during code reviews to audit resolved dependencies, and updated by automated dependency bots during dependency updates.
Table 1: The seven package managers analyzed in this study, including their associated programming languages, registries, dependency specifications, and lockfiles.
@@ 2.2 Considered Package Managers
We select seven popular package managers for our study based on the popularity of programming languages according to IEEE Spectrum in 2024 [8]. From the ten most popular languages, we identify those with widely used package managers that support lockfiles: JavaScript, TypeScript, Python, Java, and Go. We add Rust due to its significant upward trend. For each language, we select the most commonly used package manager with official support for lockfiles: npm for JavaScript/TypeScript, Poetry for Python, Cargo for Rust, Gradle for Java, and Go’s built-in toolchain. When a language has multiple package managers, we add a second one: pnpm for JavaScript and Pipenv for Python. We use GitHub as the source for identifying the most widely used package managers.Table 1 outlines the seven package managers analyzed in this study, along with their associated languages, registries, dependency specification files, and lockfiles.
@@
For instance, npm CLI is the default package manager for JavaScript, and the npm registry is the default package registry. Library developers can upload their JavaScript code to the npmjs.com registry. The dependency specification file and lockfile for npm CLI are in JSON format and are named package.json and package-lock.json, respectively. Excerpts from a package.json and its corresponding package-lock.json are given in 1 and 2, respectively. These excerpts show the basic types of metadata present in the lockfiles, such as the project name, project version, and list of dependencies. Note that the lockfile includes the exact version resolved for each dependency under the field version, along with a link to the source and its checksum under the fields resolved and integrity.
@@
In the case of Go, there is no centralized package registry. Instead, Go modules are downloaded from version control registries such as GitHub. Module versions are tracked using an index that feeds available new versions. The integrity verification of Go packages is done based on a checksum database. Moreover, in Go, the go.mod file works both as the dependency specification file and the lockfile. An additional file, named go.sum, is generated by the package manager, to record the checksums of the downloaded modules. The go.mod file is not meant to be edited manually by developers but should be modified using specific commands. 3 and 4 show an example go.mod file and a go.sum file. In this example, the go.mod file contains the module name, Go version, and the dependencies, whereas the go.sum includes the checksums of the resolved dependencies.
@@
Table 2 provides the list of selected package manager versions, including the number of lines of code in the analyzed version, the number of available releases up to that version, and the date of the first release, which together indicate the complexity and maturity of these package managers. The release counts and first release dates are obtained from the official releases page of each package manager, except for pnpm, where they are based on information from the npm registry.
@@
Table 2: Details about the exact versions of the seven package managers under study.
@@
Package manager Analyzed Version Lines of code Versions First release
npm CLI npm/cli (11.1.0) 459k (JavaScript) 557 12/2013
pnpm pnpm/pnpm (10.1.0) 112k (TypeScript) 1,103 06/2017
Cargo rust-lang/cargo (0.85.0 ) 210k (Rust) 119 10/2018
Poetry python-poetry/poetry (2.0.1) 46k (Python) 88 03/2018
Pipenv pypa/pipenv (2024.4.1) 140k (Python) 396 06/2018
Gradle gradle/gradle (8.12.1) 784k (Groovy) 531k (Java) 367k (Kotlin) 148 07/2009
Go golang/go (1.24.0) 2,052k (Go) 243 03/2012
@@ 2. RQ2:
    To what extent do projects include lockfiles in version control?
    In this RQ, we quantitatively evaluate how developers of different ecosystems include lockfiles in version control. Including lockfiles in version control systems is important to achieve all benefits lockfiles have to offer, including deterministic builds [12]. We base our evaluation on lockfile commit data collected from public GitHub projects created within the past 5 years (from 2019 to 2024).
@@ 3.2 Methodology for RQ1
To answer the RQ1, we analyze the implementations of the lockfile feature in the source code of package managers, and we review their documentation.
Source code analysis: To understand the workflows of lockfile-related tasks for each package manager, we analyze their implementation within their source code repository. We begin by locating the main function that is invoked after executing an install or build command. We then analyze the functions called by this main function, manually tracing their execution flow. We further search the project using regular expressions for the terms: lockfile, locked, frozen, and dependency resolution. We use the term dependency resolution as an umbrella term because lockfile related functions are generally tied to the dependency resolution processes. We also look for the terms locked and frozen as they are the options that users can specify to enable lockfile based functions in many package managers. The sections of source code, analyzed for each package manager are summarized in Table 3. For instance, in the case of npm, we study the repository npm/cli, focusing on the source code for install and ci actions, dependency tree resolution, and lockfile validation.
Documentation: In addition to analyzing the source code of the seven package managers, we review their official documentation. We focus on sections related to dependency resolution actions such as install or build, and study the available configurations or flags that can be used with these actions. We search for sections covering lockfiles, including lockfile-related recommendations and comparisons between lockfiles and dependency specification files. For example, as given in Table 3, for Poetry, we go through the section installing dependencies that explain lockfile usage, including installing without poetry.lock, installing with poetry.lock, and committing poetry.lock to version control. Similarly, for Pipenv, we analyze the commands and Pipfile & Pipfile.lock sections that cover general lockfile recommendations, examples, and security features. Two authors of this paper conduct this analysis, and a third one contributes in case of a disagreement. Together, they decide on the features to analyze after an initial review of all repositories and documentation. In the second run, the two authors evaluate each package manager individually against the selected features. When the two authors do not agree, they refer to the opinion of the third author. In the end, they reach full agreement, with a documented pointer to the code and documentation that describe each feature of interest.
Table 3: Lockfile features for the seven package managers under study. We study their source code and documentations.
@@ 3.3 Results for RQ1: How do package managers generate, update, and structure lockfiles?
In this RQ, we first look at the fundamental differences in the lockfile content among the seven selected package managers introduced in Table 1. Next we discuss the key differences in lockfile lifecycles including lockfile generation, updates and enforcement.